OpenStudy (anonymous):

Identify the axis of symmetry for the graph of f(x) = (x - 3)2 + 1. Enter your answer as the following example: x = 1

OpenStudy (compassionate):

Axis of symmetry: x = -b/2a

OpenStudy (compassionate):

Distribute, add like terms, and plug the terms into the equation

OpenStudy (anonymous):

how do u add like terms?

OpenStudy (callisto):

x+2x = 3x ...

OpenStudy (callisto):

Add the coefficients.

OpenStudy (anonymous):

what?

OpenStudy (callisto):



OpenStudy (anonymous):

i got f(x)=2x-5

OpenStudy (callisto):

What's that?

OpenStudy (anonymous):

idk lol

OpenStudy (callisto):

Will be back later. But f(x) = 2x-5 is not the answer to this symmetry question. I made a mistake in my previous post, the correct version is like this: If the quadratic equation is in the form of y = a(x-h)^2 +k, the axis of symmetry is x=h, the min/max value is k and the vertex is at (h,k). Note that the question given in your question is in this form, so compare it with your equation, find the value of h.